# Think Twice Before You Judge: Mixture of Dual Reasoning Experts for Multimodal Sarcasm Detection

### v1

@@ Abstract.
Multimodal sarcasm detection has attracted growing interest due to the rise of multimedia posts on social media. Understanding sarcastic image-text posts often requires external contextual knowledge, such as cultural references or commonsense reasoning. However, existing models struggle to capture the deeper rationale behind sarcasm, relying mainly on shallow cues like image captions or object-attribute pairs from images. To address this, we propose MiDRE (Mixture of Dual Reasoning Experts), which integrates an internal reasoning expert for detecting incongruities within the image-text pair and an external reasoning expert that utilizes structured rationales generated via Chain-of-Thought prompting to a Large Vision-Language Model. An adaptive gating mechanism dynamically weighs the two experts, selecting the most relevant reasoning path. Experiments on two benchmark datasets show that MiDRE achieves superior performance over baselines. Various qualitative analyses highlight the crucial role of external rationales, revealing that even when they are occasionally noisy, they provide valuable cues that guide the model toward a better understanding of sarcasm.
@@
Figure 1. Comparison between external knowledge used by existing approaches vs our approach.
@@ 1. Introduction
Sarcasm is a complex figure of speech where the intended meaning of an utterance often differs from the literal meaning. Due to its subtle and sophisticated nature, sarcasm is often employed as a tool for trolling or mocking others on social media platforms. Users intentionally exploit the gap between literal meaning and intended sentiment to convey irony, criticism, or humor in a veiled manner, making detection particularly challenging (Tindale and Gough, 1987; Averbeck, 2013; Maynard and Greenwood, 2014; Badlani et al., 2019; Ghosh et al., 2021). With the rise in multimedia content on these platforms, research towards the detection of image-text sarcasm has gained attention in recent years.
Understanding and analyzing sarcastic image-text posts poses a significant challenge, as their implicit meaning is not always conveyed directly through the surface-level text or visual content, frequently relying on external contextual information rather than just the content itself. For instance, in Figure 1(a), the caption claims a shortage of cabs, while the image clearly depicts an abundance of them. This form of sarcasm is direct and easily recognizable without requiring external knowledge. In Figure 1(b), the caption ”time to change the channel for some good news” mocks the idea of escaping a storm warning by simply switching channels. While still indirect, the sarcasm is relatively easy to grasp through description of visual content. However, in Figure 1(c), which suggests Halloween items might appear as early as March or April, relies on knowledge of seasonal retail patterns to detect the sarcasm, which makes it extremely difficult to detect.
@@
Hence, we argue the need for rich contextual reasoning, where external knowledge is not only verbose but also structured into modality-specific reasoning, for gaining multiple viewpoints from the input. Unlike generic reasoning as used by earlier models, these reasonings should be sarcasm-oriented, offering clear and relevant cues. In designing our approach, we are guided by the following key observations:
@@
    Hallucination and bias in generated rationales: While LVLM-generated rationales can be informative, they may include hallucinations or factual errors. Relying on them blindly can mislead the model. Therefore, our approach should combine these external cues with cues from the input image-text for more reliable predictions.
To this end, we propose MiDRE (Mixture of Dual Reasoning Experts), a model designed to explicitly combine external reasoning cues with internal reasoning from input image-text for multimodal sarcasm detection. To realise observations (1) and (2), we generate external contextual knowledge from a pretrained Large Vision-Language Model (LVLM) using chain-of-thought (CoT) prompting, guiding it to produce step-by-step rationales: first from the image, then the text, and finally from the combined image-text input.
To alleviate the hallucination problem specified in observation (3), MiDRE introduces two complementary expert networks (internal expert and external expert) within an LLM backbone. The internal expert considers the raw image-text pair, capturing sarcasm from surface-level incongruities, without relying on external information. To capture sarcastic cues from external knowledge, external expert considers the image-text content as well as the rationales generated from the LVLM. A learnable gating module dynamically fuses the outputs from both experts, enabling MiDRE to adaptively balance between noisy external rationales and cues from the input image-text. This selective integration leads to more robust and interpretable sarcasm detection. Thus our key contributions are as follows:
@@
  • To address the semantic gap of contextual knowledge, we leverage multi-level external rationales (image, text, and combined) generated by a pretrained LVLM with chain-of-thought prompting for sarcasm detection.
  • We conduct extensive experiments and demonstrate that MiDRE outperforms existing baselines, providing both improved accuracy and interpretability in sarcasm detection.
@@ 2.1. Image-Text Sarcasm Detection
Figure 2. Rationale generation module of MiDRE. Detailed prompts are presented in appendix section A.1.
@@ 2.2. Mixture of Experts
In contrast to existing approaches, we propose a novel framework that addresses the problem of sarcasm detection (image+text) from dual perspective, content and external rationales. Our dual expert approach can identify saracsm based on considering the content or external rationales or a mix of both.
@@ 3.1. Task Definition
The task of image-text sarcasm detection can be framed as a binary classification problem where, given a multimodal input sample pair $x=(T,I)$, where $T$ is the textual content and $I$ is the image, the objective is to assign $x$ a label $y$ from the set $Y=\{yes,no\}$ where $yes$ means sarcastic.
@@ 3.2. Architecture of MiDRE
Our proposed model, MiDRE, detects sarcasm by reasoning from two complementary perspectives: (1) the content itself, and (2) the content enriched with external rationales. As shown in Figure 3, the internal reasoning expert ($IR$) captures inherent sarcastic cues directly from the input, using visual and textual features extracted via CLIP encoders. To identify sarcasm using both content and external knowledge, we first introduce a Rationale Generation Module that leverages a pre-trained LVLM with chain-of-thought prompting to generate structured, modality-specific sarcasm rationales across three views: image, text, and combined image-text. The external reasoning expert ($ER$) performs deeper reasoning using LLM-encoded representations of the input text and generated rationales, alongside CLIP-based image embeddings for visual context. A learnable gating network then adaptively fuses the outputs of $ER$ and the $IR$, allowing the model to balance reliance on external rationales versus direct image-text cues for more accurate and robust sarcasm detection. This combination of $IR$, $ER$ and gating is incorporated in all encoder layers of an encoder-decoder based LLM.
@@ 3.3. Rationale Generation Module
We observe that when the LVLM was prompted with both the image and text to generate all three rationales together, it often blended the sources—such as incorporating image details into the text-only rationale, and vice versa. To avoid this cross-modal interference and ensure that each rationale reflects its intended modality, we designed an independent three-stage extraction process. This setup helps maintain clear boundaries between image, text, and combined reasoning.
The three-stage process is described below.
Stage 1- Image-based rationale generation: In this stage, the LVLM is prompted to focus exclusively on the image, instructing it to describe the image content, infer possible implications, and reason whether the visual content alone could indicate sarcasm or irony. This helps capture cues like meme-style content or ironic visual contexts. This stage can be formulated as: $R^{I}=LVLM(Prompt^{I},I)$
@@
We combine the three rationales to form a comprehensive external reasoning input $R=[R^{I},R^{T},R^{M}]$. This combined rationale encapsulates detailed external reasoning signals from each modality and their interaction. The order of combination of rationales is arbitrary as we observed very little variation in performance during our experiments.
@@ 3.4. Dual Reasoning Expert Networks
The internal reasoning expert $(IR)$, external reasoning expert $(ER)$ and the adaptive gate are deeply integrated into every encoder layers of the LLM. Once the rationale $R$ is extracted from the rationale generation module, the modified input sample becomes $x=\{I,T,\hat{T}\}$, where $\hat{T}=[T,R]$ is the combination of text and rationale. $ER$ considers the image-text along with rationales as input. The input required for $ER$ is extracted as
(1) $H_{I}=VE(I),\quad H_{\hat{T}}^{0}=E_{LLM}(\hat{T})$
where $H_{I}\in\mathbb{R}^{m\times d_{v}}$ is the image embedding from CLIP visual encoder $VE(\cdot)$, and $m$ is the number of image patches. $H_{\hat{T}}^{0}\in\mathbb{R}^{n\times d_{t}}$ is the embedding of the combined text and rationale input $\hat{T}$ from LLM embedding layer $E_{LLM}(\cdot)$, where $n$ is the token length. Since image embeddings $H_{I}$ is from the CLIP image encoder, to align it with the LLM embedding space, we use a separate learnable projection layer:
(2) $H_{I^{\prime}}=H_{I}W_{i}$
@@
$\hat{T}$ contains richer and more complex semantic information than $T$, hence we pass it through the transformer layers of the encoder to leverage the deep contextual reasoning capabilities of the LLM and facilitate more sophisticated integration of external rationale signals.
Specifically in LLM encoder layer ${i}$, the computation of $ER$ can be formulated as:
(3) $M_{ER}^{i}=ER^{i}(H_{\hat{T}}^{i-1},H_{I^{\prime}})$
where $M_{ER}^{i}$ is the output from the external reasoning expert $ER^{i}$ in the encoder layer $i$. $ER$ is implemented like a cross-attention mechanism as follows:
@@
(4) $ER(H_{\hat{T}},H_{I^{\prime}})=\text{softmax}\left(\frac{Q_{\hat{T}}K^{T}_{I^{\prime}}}{\sqrt{d}}\right)V_{I^{\prime}}$
where $\{Q_{\hat{T}},K^{T}_{I^{\prime}},V_{I^{\prime}}\}=\{H_{\hat{T}}W_{Q},H_{I^{\prime}}W_{K},H_{I^{\prime}}W_{V}\}$. $\{W_{Q},W_{K},W_{V}\}\in\mathbb{R}^{d\times d_{k}}$ are query, key and value projections respectively.
The internal reasoning expert ($IR$) is designed to learn semantic associations directly from the input image-text pair without relying on any external rationales. Since $T$ is comparatively simpler than the enriched rationale-augmented input $\hat{T}$, and to avoid potential interference or leakage of rationale information into the internal reasoning stream, we extract the embedding of $T$ separately using a CLIP text encoder. This design choice ensures that $IR$ operates purely on the original image and text content. To validate this design choice, we also experimented with a reversed configuration—feeding $T$ into the LLM encoder and $\hat{T}$ into the CLIP text encoder, but this led to inferior performance, as reported in Appendix A.2.
The text input required for $IR$ is extracted as $H_{T}=VE(T)$
where $H_{T}\in\mathbb{R}^{n\times d_{t}}$ is the embedding of text only input $T$ from CLIP text encoder $TE(\cdot)$. Since text embeddings $H_{T}$ is from CLIP encoder, to align it with the LLM embedding space, we use a separate learnable projection layer:
(5) $H_{T^{\prime}}=H_{T}W_{t}$
We utilize the same image embeddings ${H}_{I^{\prime}}$ as of $ER$. At each LLM encoder layer $i$, it produces the output:
(6) $M_{IR}^{i}=IR^{i}(H_{T^{\prime}},H_{I^{\prime}})$
The internal reasoning expert is realized as:
(7) $IR(H_{T^{\prime}},H_{I^{\prime}})=\text{softmax}\left(\frac{Q_{T^{\prime}}K^{T}_{I^{\prime}}}{\sqrt{d}}\right)V_{I^{\prime}}$
where $\{Q_{T^{\prime}},K_{I^{\prime}},V_{I^{\prime}}\}=\{H_{T^{\prime}}W_{Q},H_{I^{\prime}}W_{K},H_{I^{\prime}}W_{V}\}$.
While $IR$ suffices in cases where the text and image directly convey sarcastic cues, there are scenarios where IR alone is inadequate, and the model must rely on the $ER$, for contextual understanding. However, external reasoning can also be noisy or unreliable. To adaptively balance these signals, we introduce a gating mechanism that dynamically weighs the contributions of $IR$ and $ER$ at each encoder layer $i$. The gating is realized as a bottleneck layer that takes the concatenated embeddings of both expert outputs:
(8) $G^{i}=\text{softmax}(W_{g2}^{i}\cdot\text{GELU}\left(W_{g1}^{i}[M_{ER}^{i};M_{IR}^{i}]\right))$
where $W_{g1}^{i}\in\mathbb{R}^{2d_{t}\times d_{b}}$, $W_{g2}^{i}\in\mathbb{R}^{d_{b}\times 2}$, $d_{b}<d$ defines the bottleneck dimension, and $G^{i}\in\mathbb{R}^{2}$.
The final gated output is then computed as:
(9) $M_{G}^{i}=G^{i}_{1}\odot M_{ER}^{i}\;+\;G^{i}_{2}\odot M_{IR}^{i}$
The bottleneck design allows the gate to act as an information filter to dynamically attend to either $IR$, $ER$, or a context-dependent combination of both. We utilize a residual connection to generate the final output from encoder layer $i$ which will be fed to the next encoder layer $i+1$:
(10) $H_{\hat{T}}^{i}=H_{\hat{T}}^{i-1}+M_{G}^{i}$

### v2

@@ Abstract.
Multimodal sarcasm detection has attracted growing interest due to the rise of multimedia posts on social media. Understanding sarcastic image-text posts often requires external contextual knowledge, such as cultural references or commonsense reasoning. However, existing models struggle to capture the deeper rationale behind sarcasm, relying mainly on shallow cues like image captions or object-attribute pairs from images. To address this, we propose MiDRE (Mixture of Dual Reasoning Experts), which integrates an internal reasoning expert for detecting incongruities within the image-text pair and an external reasoning expert that utilizes structured rationales generated via Chain-of-Thought prompting to a Large Vision-Language Model. An adaptive gating mechanism dynamically weighs the two experts, selecting the most relevant reasoning path. Unlike prior methods that treat external knowledge as static input, MiDRE selectively adapts to when such knowledge is beneficial, mitigating the risks of hallucinated or irrelevant signals from large models. Experiments on two benchmark datasets show that MiDRE achieves superior performance over baselines. Various qualitative analyses highlight the crucial role of external rationales, revealing that even when they are occasionally noisy, they provide valuable cues that guide the model toward a better understanding of sarcasm.
@@
Figure 1. Comparison between external knowledge used by existing approaches vs our approach.:
@@ 1. Introduction
Sarcasm is a complex figure of speech where the intended meaning of an utterance often differs from the literal meaning. Due to its subtle and sophisticated nature, sarcasm is often employed as a tool for trolling or mocking others on social media platforms. Users intentionally exploit the gap between literal meaning and intended sentiment to convey irony, criticism, or humor in a veiled manner, making detection particularly challenging (Tindale and Gough, 1987; Averbeck, 2013; Maynard and Greenwood, 2014; Badlani et al., 2019; Ghosh et al., 2021). With the rise in multimedia content on these platforms, research towards the detection of image-text sarcasm has gained significant attention in recent years.
Understanding and analyzing sarcastic image-text posts poses a significant challenge, as their implicit meaning is not always conveyed directly through the surface-level text or visual content, frequently relying on external contextual information. For instance, in Figure 1(a), the caption claims a shortage of cabs, while the image clearly depicts an abundance of them. This form of sarcasm is direct and easily recognizable without requiring external knowledge. In Figure 1(b), the caption ”time to change the channel for some good news” mocks the idea of escaping a storm warning by simply switching channels. While still indirect, the sarcasm is relatively easy to grasp through description of visual content. However, in Figure 1(c), which suggests Halloween items might appear as early as March or April, relies on knowledge of seasonal retail patterns to detect the sarcasm, which makes it extremely difficult to detect.
@@
Hence, we argue the need for rich contextual reasoning, where external knowledge is not only verbose but also structured into modality-specific reasoning, for gaining multiple viewpoints from the input. Unlike generic external knowledge used by earlier models, these reasonings should be sarcasm-oriented, offering clear and relevant cues. In designing our approach, we are guided by the following key observations:
@@
    Hallucination and bias in generated rationales: While LVLM-generated rationales can be informative, they may include hallucinations or factual errors. Relying on them blindly can mislead the model. Therefore, our approach should combine these rationales with cues from the input image-text for more reliable predictions.
To this end, we propose MiDRE (Mixture of Dual Reasoning Experts), a model designed to explicitly combine external reasoning with internal reasoning from input image-text for multimodal sarcasm detection. To realise observations (1) and (2), we generate rationales from a LVLM using chain-of-thought (CoT) prompting, guiding it to produce step-by-step rationales: first from the image, then the text, and finally from the combined image-text input.
To alleviate the hallucination problem specified in observation (3), MiDRE introduces two complementary expert networks: an internal reasoning expert network (IRE) and an external reasoning expert network (ERE) within an LLM backbone. The IRE considers the raw image-text pair, capturing sarcasm from surface-level incongruities, without relying on external information. To capture sarcastic cues from external knowledge, ERE considers the image-text content as well as the rationales generated from the LVLM. A learnable gating module dynamically fuses the outputs from both experts, enabling MiDRE to adaptively balance between noisy external rationales and cues from the input image-text. This selective integration leads to more robust and interpretable sarcasm detection. MiDRE achieves notable improvements over the best-performing baselines, with accuracy gains of $\Delta$1.84 and $\Delta$1.04 and F1-score gains of $\Delta$1.02 and $\Delta$3.79 on MMSD and MMSD2.0, respectively. Our key contributions are as follows:
@@
  • To address the semantic gap of contextual knowledge, we leverage multi-level external rationales (image, text, and combined) generated by a pretrained LVLM with chain-of-thought prompting.
  • Through extensive experimentation, we show that MiDRE consistently surpasses existing baselines, achieving higher performance while making it possible to trace and understand the model’s reasoning pathway via gating patterns.
@@ 2.1. Image-Text Sarcasm Detection
Figure 2. Rationale generation module of MiDRE.
@@ 2.2. Mixture of Experts
In contrast, we propose a novel framework that addresses the problem of sarcasm detection (image+text) from dual perspective, content and external rationales. Our dual expert approach can identify saracsm based on considering the content or external rationales or a mix of both.
@@ 3.1. Task Definition
The task of image-text sarcasm detection can be framed as a binary classification problem where, given a multimodal input sample pair $x=(T,I)$, where $T$ is the textual content and $I$ is the image, the objective is to assign $x$ a label $y$ from the set $Y=\{yes,no\}$ where yesyes means sarcastic.
@@ 3.2. Architecture of MiDRE
Our proposed model, MiDRE, detects sarcasm by reasoning from two complementary perspectives: (1) the content itself, and (2) the content enriched with external rationales. As shown in Figure 3, the internal reasoning expert IRE captures inherent sarcastic cues directly from the input, using visual and textual features extracted via CLIP encoders. To identify sarcasm using both content and external knowledge, we first introduce a Rationale Generation Module that leverages a pre-trained LVLM with CoT prompting to generate structured, modality-specific sarcasm rationales across three views: image, text, and combined image-text. The external reasoning expert ERE performs deeper reasoning using LLM-embedded representations of the input text and generated rationales, alongside CLIP-based image embeddings. A learnable gating network then adaptively fuses the outputs of ERE and the IRE, allowing the model to balance its reliance on external rationales versus direct image-text cues for more accurate and robust sarcasm detection. This combination of IRE, ERE and gating is incorporated in all encoder layers of an encoder-decoder based LLM to enable dynamic expert selection at multiple levels of representation. This would ensure that both internal and external reasoning cues are effectively leveraged throughout the model’s processing pipeline.
@@ 3.3. Rationale Generation Module
We observe that when the LVLM was prompted with both the image and text to generate all three rationales together, it often blended the sources—such as incorporating image details into the text-only rationale, and vice versa. To avoid this cross-modal interference and ensure that each rationale reflects its intended modality, we designed an independent three-stage extraction process. This setup helps maintain clear boundaries between image, text, and combined reasoning. The three-stage process is described below.
Stage 1- Image-based rationale generation: In this stage, the LVLM is prompted to focus exclusively on the image, instructing it to describe the image content and reason whether the visual content indicates sarcasm. This helps capture cues like meme-style content or ironic visual contexts. This stage can be formulated as: $R^{I}=LVLM(Prompt^{I},I)$
@@
We combine the three rationales to form a comprehensive external reasoning input $R=[R^{I},R^{T},R^{M}]$. This combined rationale encapsulates detailed external reasoning signals from each modality and their interaction. The order of combination is arbitrary, as we observed very little variation in performance during our experiments.
@@ 3.4. Dual Reasoning Expert Networks
Our model integrates the Internal Reasoning Expert (IRE), the External Reasoning Expert (ERE), and an adaptive gating mechanism into every encoder layer of the LLM. This layered design enables the model to dynamically combine reasoning signals derived from both original inputs and auxiliary rationales.
External Reasoning Expert (ERE): Once the rationale $R$ is extracted from the rationale generation module, we construct a rationale-augmented input sample
$x=\{I,T,P\},\quad P=[T,R],$
where $I$ is the image, $T$ is the original text, and $P$ is the concatenation of text and rationale.
The $ERE$ takes as input the image $I$ and the enriched sequence $P$. Image features are extracted using a CLIP visual encoder:
$H_{I}=VE(I),\quad H_{I}\in\mathbb{R}^{m\times d_{v}},$
@@
where $m$ is the number of image patches. The sequence $P$ contains richer semantic information than $T$ alone, so we pass it through the LLM’s transformer encoder layers to capture deeper contextual dependencies and enable sophisticated integration of external rationale signals. $P$ is embedded by the LLM embedding layer:
$H_{P}^{0}=Emb_{\text{LLM}}(P),\quad H_{P}^{0}\in\mathbb{R}^{n\times d_{t}},$
where $n$ is the token length of $P$, $d_{t}$ is the LLM embedding dimension, and $H_{P}^{0}$ represents the initial embedding from the LLM that will be fed to the first encoder layer. To align the CLIP visual features with the LLM’s embedding space, we apply a learnable projection:
$H_{I^{\prime}}=H_{I}W_{i}.$
In encoder layer $i$, the $ERE$ output is computed as:
$M_{ERE}^{i}=ERE^{i}(H_{P}^{i-1},H_{I^{\prime}}),$
where $ERE$ is implemented as a cross-attention mechanism:
$ERE(H_{P},H_{I^{\prime}})=\text{softmax}\!\left(\frac{Q_{P}K_{I^{\prime}}^{T}}{\sqrt{d}}\right)V_{I^{\prime}},$
with $\{Q_{P},K_{I^{\prime}},V_{I^{\prime}}\}=\{H_{P}W_{Q},\,H_{I^{\prime}}W_{K},\,H_{I^{\prime}}W_{V}\}$ and $\{W_{Q},W_{K},W_{V}\}\in\mathbb{R}^{d\times d_{k}}$.
Internal Reasoning Expert (IRE): While the $ERE$ leverages rationale-augmented inputs, the $IRE$ is designed to operate only on the original image-text pair $(I,T)$, learning semantic associations without external rationale guidance. Since $T$ is a simpler sequence compared to the enriched input $P$, passing it through the LLM encoder is unnecessary and could introduce interference with $P$’s processing. Instead, $T$ is encoded using the CLIP text encoder, which is sufficient to capture its semantic content while keeping the LLM encoder dedicated to the more complex reasoning required for $P$.
We embed $T$ using the CLIP text encoder and project it to the LLM embedding space to get $H_{T^{\prime}}$:
$H_{T}=TE(T),\quad H_{T}\in\mathbb{R}^{n\times d_{t}},\quad H_{T^{\prime}}=H_{T}W_{t}.$
The $IRE$ uses the same projected image features $H_{I^{\prime}}$ as the $ERE$, and in layer $i$ computes:
$M_{IRE}^{i}=IRE^{i}(H_{T^{\prime}},H_{I^{\prime}}),$
@@
$\text{where,}\hskip 5.69046ptIRE(H_{T^{\prime}},H_{I^{\prime}})=\text{softmax}\!\left(\frac{Q_{T^{\prime}}K_{I^{\prime}}^{T}}{\sqrt{d}}\right)V_{I^{\prime}},$
and $\{Q_{T^{\prime}},K_{I^{\prime}},V_{I^{\prime}}\}=\{H_{T^{\prime}}W_{Q},\,H_{I^{\prime}}W_{K},\,H_{I^{\prime}}W_{V}\}$.
Design Choice: $P$ is processed by the LLM encoder and $T$ by the CLIP text encoder, reflecting the higher complexity of $P$ relative to $T$. Reversing this configuration, assigning $T$ to the LLM encoder and $P$ to CLIP text encoder resulted in lower performance, reported in Table 5 (row 7) indicated by w Swap Enc (P $\leftrightarrow$ T).
Adaptive Gating: In many cases, $IRE$ alone suffices for sarcasm detection when the image and text provide strong intrinsic cues. However, certain examples require contextual knowledge captured by $ERE$, while in other cases, $ERE$ may introduce noise. To dynamically balance these contributions, we employ a trainable gating mechanism in each encoder layer $i$.
The gate takes the concatenated outputs of both experts:
$G^{i}=\text{softmax}\!\left(W_{g2}^{i}\cdot\text{GELU}\left(W_{g1}^{i}[M_{ERE}^{i};\,M_{IRE}^{i}]\right)\right),$
where $W_{g1}^{i}\in\mathbb{R}^{2d_{t}\times d_{b}}$, $W_{g2}^{i}\in\mathbb{R}^{d_{b}\times 2}$, $d_{b}<d$ is the bottleneck dimension, and $G^{i}\in\mathbb{R}^{2}$.
The gated output is then computed as:
$M_{G}^{i}=\big(G^{i}[1]\big)\odot M_{ERE}^{i}\;+\;\big(G^{i}[2]\big)\odot M_{IRE}^{i},$
We apply a residual connection to obtain the final representation for the next encoder layer:
$H_{P}^{i}=H_{P}^{i-1}+M_{G}^{i}.$
This design allows the model to selectively attend to internal reasoning, external reasoning, or a context-dependent combination of both at each layer, ensuring robustness to noisy rationales while retaining the benefits of enriched contextual cues.